OpenStudy (theraggedydoctor):

On a blueprint, 0.4 inch represents 6 feet. If the actual distance between two buildings is 76 feet, what would be the distance between the corresponding buildings on the blueprint?

OpenStudy (theraggedydoctor):

\(\Large\frac{ 0.4 }{ ? }=\frac{ 6 }{ 76 }\) I multiplied \(76 \times \ 0.4\) and then divided \(30.4 \div \ 6\) and I got \(5.06...7\) and rounded that to \(5.1\).

OpenStudy (theraggedydoctor):

Is that right?!

OpenStudy (anonymous):

Give me a sec... 0.4 in : 6 ft this is the scale ?? : 76 ft it should be like this